OpenStudy (anonymous):

how i can solve this x+9=5;4

OpenStudy (anonymous):

what does the ";" mean?

OpenStudy (anonymous):

-4 +9 = 5

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

on my paper says check wheather the given number is a solution of the equation

OpenStudy (gw2011):

x+9=5 x=5-9 x=-4 Therefore, 4 is not a solution to this equation but -4 is the solution